OpenStudy (anonymous):

(x-3)(x^2+2x+4) mutiply. Please show work

OpenStudy (anonymous):

just distribute (x-3)(x^2+2x+4=x^3+2x^2+4x-3x^2-6x-12=x^3-x^2-2x-12

OpenStudy (anonymous):